# Minibatch Optimal Transport and Perplexity Bound Estimation in Discrete Flow Matching

### v3

@@ 5.4 Scaling Properties of Minibatch-OT
We examine the computational overhead introduced by minibatch OT during training. While OT computation is vocabulary-size independent, its requirements increase with batch size under the Sinkhorn/exact algorithms and each parameter update requires computing a minibatch OT coupling. Table 4 compares the time for 1000 couplings (CPU or GPU) against 1000 diffusion updates without OT. OT adds only 3.4% overhead in our experiments. Larger batch sizes require GPU acceleration, maintaining the overhead between 10-15%. All experiments use fixed sequence length $L=128$.
@@
Pure diffusion 54.7 63.6 104.9 173.0 367.8 634.4 $1129^{E}$ $2010^{E}$
Sequence length does not have a negative impact on computational scaling. Normally, the primary overhead stems from the Sinkhorn operation, which processes pre-computed pairwise sequence distances. Consequently, sequence length does not affect Sinkhorn’s computational cost. We tested the overall role of the length empirically by increasing the sequence length 8 times (from $L=128$ to $L=1024$), which yielded only a 4.6x increase in OT computation time (from 12,57 to 57,99 seconds per thousand minibatches). This scaling behavior has important implications for training efficiency: At $L=128$, OT adds 3.4% to the total training time. At $L=1024$, this overhead should drop below 1.9% because diffusion-only training time scales at best roughly linearly (and up to quadratically if attention dominates) with sequence length, whereas OT in our experiments scaled more slowly. Consequently, the relative cost of mini-batch OT is not expected to increase with sequence length.
In addition, minibatch OT scales favorably with embedding dimension $D$: OT computation scales linearly in embedding dimension $D$, while transformer computation scales quadratically in $D$. Consequently, as the MLPs and embedding dimensions in the transformer grow, the relative overhead of our proposed approach decreases compared to the flow optimization cost. Furthermore, our method is agnostic to the number of transformer layers, which means the relative overhead is expected to decrease even further as models scale.
5.5 Ablation Experiments
We conduct ablation studies on two key factors: OT batch size and OT solver. Switching from Sinkhorn to exact solver in multimask flow yields minor improvements for small step sizes (8 and 16), with more pronounced gains in bound values. Increasing minibatch OT size to 4096 (while maintaining diffusion batch size) slightly improves results at 32 and 64 steps, again with more pronounced improvements in bound values. More details and the results can be found in Appendix C.8.
@@ 6 Related Work
Diffusion-based models have proven highly effective in capturing the structure of continuous data distributions, leading to significant advancements in generative modeling (Song et al., 2020a; 2021c; Kingma et al., 2021; Nichol & Dhariwal, 2021; Saharia et al., 2022; Ramesh et al., 2022). Given their success in image, video and audio synthesis, researchers have explored their applicability to language modeling (Chen et al., 2023; Gulrajani & Hashimoto, 2024; Li et al., 2022; Dieleman et al., 2022; Strudel et al., 2022; Gong et al., 2022; Mahabadi et al., 2023).
An alternative paradigm for discrete data, particularly in NLP, is discrete diffusion. Introduced by Hoogeboom et al. (2021); Austin et al. (2021) and extended to continuous-time settings (Campbell et al., 2022; Lou et al., 2024), these models offer a structured approach to learning categorical distributions. Training typically uses the variational lower bound or cross-entropy loss, similar to continuous diffusion (Dieleman et al., 2022).
To expand the design space of discrete diffusion, Campbell et al. (2024); Gat et al. (2024) introduce discrete flow matching, notably avoiding conditional score ratio calculations during training and thus bypassing matrix exponential computation. Instead of focusing on a path-length-oriented objective, Shaul et al. (2025) define a kinetic energy OT objective, derive the optimum for specific DFM classes, and independently obtain a bound similar to ours from an ELBO perspective. While in this work we focus on pure DFM models, Arriola et al. (2025) introduce block diffusion language models that interpolate between discrete denoising diffusion and autoregressive models. Regarding scaling, Nie et al. (2025) train masked diffusion models up to 1.1B parameters to systematically evaluate against comparable or larger ARMs. Their 1.1B MDM outperforms the 1.1B TinyLlama trained on the same data across four of eight zero-shot benchmarks.
7 Limitations and Future Work
Flow matching with minibatch OT involves two interacting optimization procedures: choosing optimal minibatch coupling and training the flow model. The coupling affects flow dynamics, while embedding updates during training alter the optimal coupling when using embedding-based similarities. Though the model’s local view of the flow at each step provides stability, we can enhance it by decoupling network embeddings from those used in minibatch coupling, for instance, using moving-average embeddings for OT.
@@ 8 Conclusion
We developed a weighted path-length dynamic OT objective for DFM that minimizes dissimilarity-weighted jumps between states, derived its Kantorovich formulation establishing a categorical Benamou-Brenier-type theorem. We extended two discrete diffusion bounds to DFM, enabling comparisons with autoregressive and discrete diffusion models. Experiments show minibatch OT reduces inference steps up to 8-fold (1024 to 128) while maintaining generative perplexity on GPT2-scale models. Our multimask flow (DFM-MMLM) surpasses masked DFM in generative perplexity without sacrificing diversity, with further gains under OT.
LLM Usage Statement
Large Language Models were used in this paper to improve the conciseness and quality of the text at the sentence level.
APPENDIX
@@ A.1 Proofs
is clearly $s(x_{1}^{i},x_{0}^{i})$ when $x_{0}^{i}\neq x_{1}^{i}$ and zero otherwise. Thus, we have
@@
$D_{KL}(\hat{q},\hat{p})=\sum_{x_{0:K\epsilon}}\hat{q}(x_{0:K\epsilon})\log{\frac{\hat{q}(x_{0:K\epsilon})}{\hat{p}(x_{0:K\epsilon})}}=\sum_{x_{0:K\epsilon}}\hat{q}(x_{0:K\epsilon})\log{\prod_{k=1}^{K}\frac{\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon},...,x_{0})}{\hat{p}(x_{k\epsilon}|x_{(k-1)\epsilon},...,x_{0})}\frac{\hat{q}(x_{0})}{\hat{p}(x_{0})}}$ (36)
@@
First we will simplify notation and write $t_{k}=(k-1)\epsilon$, as well as $\hat{q}(x_{k\epsilon}=x|x_{(k-1)\epsilon}=z)=\hat{q}_{t_{k}+\epsilon|t_{k}}(x|z)$, where $z$ and $x$ are states. Therefore the previous Expression (41) becomes
@@
$\sum_{\begin{subarray}{c}x\\ d_{H}(x,z)=0\end{subarray}}\hat{q}_{t_{k}+\epsilon|t_{k}}(x|z)\log{\frac{\hat{q}_{t_{k}+\epsilon|t_{k}}(x|z)}{\hat{p}_{t_{k}+\epsilon|t_{k}}(x|z)}}+\sum_{\begin{subarray}{c}x\\ d_{H}(x,z)=1\end{subarray}}\hat{q}_{t_{k}+\epsilon|t_{k}}(x|z)\log{\frac{\hat{q}_{t_{k}+\epsilon|t_{k}}(x|z)}{\hat{p}_{t_{k}+\epsilon|t_{k}}(x|z)}}$ (45)
@@
$\sum_{\begin{subarray}{c}x\\ d_{H}(x,z)=1\end{subarray}}\hat{q}_{t_{k}+\epsilon|t_{k}}(x|z)\log{\frac{\hat{q}_{t_{k}+\epsilon|t_{k}}(x|z)}{\hat{p}_{t_{k}+\epsilon|t_{k}}(x|z)}}$ (48)
$\sum_{j=1}^{L}\sum_{x^{j}\neq z^{j}}\epsilon w_{t_{k}}^{j}(x^{j},z)\log{\frac{w_{t_{k}}^{j}(x^{j},z)+O(\epsilon)}{v_{t_{k}}^{j}(x^{j},z)+O(\epsilon)}}+O(\epsilon^{2}).$ (49)
@@
$D_{KL}(\bar{q},\bar{p})=D_{KL}(\bar{q}_{0}\|\bar{p}_{0})$
$\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\left(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-w_{t}^{i}(x^{i},x_{t})\right)dt.$ (63)
@@
$D_{KL}(\bar{q},\bar{p})=D_{KL}(\bar{q}_{1}\|\bar{p}_{1})$
@@
$D_{KL}(\bar{q}_{1}\|\bar{p}_{1})=D_{KL}(\bar{q}_{0}\|\bar{p}_{0})+$
$\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-w_{t}^{i}(x^{i},x_{t})\bigg)dt$-
@@
$D_{KL}(\bar{q},\bar{p})=$
$\int_{0}^{1}\sum_{z}\bar{q}_{1-t}(z)\sum_{i=1}^{L}\sum_{x^{i}\neq z^{i}}\bigg(r_{\bar{q}_{1-t}}w_{1-t}^{i}(z^{i},x)\log{\frac{r_{\bar{q}_{1-t}}w_{1-t}^{i}(z^{i},x)}{r_{\bar{p}_{1-t}}v_{1-t}^{i}(z^{i},x)}}$
$r_{\bar{p}_{1-t}}v_{1-t}^{i}(z^{i},x)-r_{\bar{q}_{1-t}}w_{1-t}^{i}(z^{i},x)\bigg)dt$
$D_{KL}(\bar{q_{1}},\bar{p_{1}}).$ (71)
@@
$D_{KL}(\bar{q},\bar{p})=D_{KL}(\bar{q_{1}},\bar{p_{1}})$
$\int_{0}^{1}\sum_{z}\bar{q}_{\tau}(z)\sum_{i=1}^{L}\sum_{x^{i}\neq z^{i}}\left(r_{\bar{q}_{\tau}}w_{\tau}^{i}(z^{i},x)\log{\frac{r_{\bar{q}_{\tau}}w_{\tau}^{i}(z^{i},x)}{r_{\bar{p}_{\tau}}v_{\tau}^{i}(z^{i},x)}}+r_{\bar{p}_{\tau}}v_{\tau}^{i}(z^{i},x)-r_{\bar{q}_{\tau}}w_{\tau}^{i}(z^{i},x)\right)d\tau$ (72)
@@
We now set to prove that $D_{KL}(\bar{q}(x_{1})\|\bar{p}(x_{1}))\leq D_{KL}(\bar{q},\bar{p})$. Since in Equation (66), the term
@@
$D_{KL}(\bar{q_{1}}\|\bar{p_{1}})\leq D_{KL}(\bar{q_{0}}\|\bar{p_{0}})$
$\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\left(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-w_{t}^{i}(x^{i},x_{t})\right)dt$ (74)
@@
$D_{KL}(\bar{q}_{1}\|\bar{p}_{1})\leq D_{KL}(\bar{q}_{0}\|\bar{p}_{0})$
$\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\left(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-w_{t}^{i}(x^{i},x_{t})\right)dt$ (78)
@@
$\frac{\partial H(\bar{q}_{t})}{\partial t}=-\frac{\partial}{\partial t}\sum_{x_{t}}\bar{q}_{t}(x_{t})\log{\bar{q}_{t}(x_{t})}=-\sum_{x_{t}}\frac{\partial\bar{q}_{t}(x_{t})}{\partial t}\left(\log{\bar{q}_{t}(x_{t})}+1\right)$ (80)
$\sum_{x_{t}}\frac{\partial\bar{q}_{t}(x_{t})}{\partial t}\left(\log{\frac{\bar{q}_{t}(x)}{\bar{q}_{t}(x_{t})}}-1\right)-\sum_{x_{t}}\frac{\partial\bar{q}_{t}(x_{t})}{\partial t}\log{\bar{q}_{t}(x)}.$ (81)
We prove the last term $\sum_{x_{t}}\frac{\partial\bar{q}_{t}(x_{t})}{\partial t}\log{\bar{q}_{t}(x)}dx_{t}$ is 0. From the Continuity Equation (Gat et al., 2024),
@@
$\sum_{x_{t}}\frac{\partial\bar{q}_{t}(x_{t})}{\partial t}\log{\bar{q}_{t}(x)}$
$\sum_{x}\bar{q}_{t}(x)\sum_{i=1}^{L}\sum_{x_{t}}\left(\delta_{x}(x_{t}^{-i})w_{t}^{i}(x_{t}^{i},x)\right)\log{\bar{q}_{t}(x)}=\sum_{x}\bar{q}_{t}(x)\sum_{i=1}^{L}0\log{\bar{q}_{t}(x)=0.}$ (83)
This implies that
$\frac{\partial H(\bar{q}_{t})}{\partial t}=\sum_{x_{t}}\frac{\partial\bar{q}_{t}(x_{t})}{\partial t}\left(\log{\frac{\bar{q}_{t}(x)}{\bar{q}_{t}(x_{t})}}-1\right)=\sum_{x_{t}}\sum_{x}\bar{q}_{t}(x)\sum_{i=1}^{L}\delta_{x}(x_{t}^{-i})w_{t}^{i}(x_{t}^{i},x)\left(\log{\frac{\bar{q}_{t}(x)}{\bar{q}_{t}(x_{t})}}-1\right)$ (84)
$\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}}w_{t}^{i}(x_{t}^{i},x)\frac{\bar{q}_{t}(x)}{\bar{q}_{t}(x_{t})}\left(\log{\frac{\bar{q}_{t}(x)}{\bar{q}_{t}(x_{t})}}-1\right).$ (85)
@@
$D_{KL}(\bar{q}_{1}\|\bar{p}_{1})\leq D_{KL}(\bar{q}_{0}\|\bar{p}_{0})+$ (87)
$\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\left(r_{\bar{q}_{t}}\tilde{w}_{t}^{i}(x_{t}^{i},x)\log{\frac{r_{\bar{q}_{t}}\tilde{w}_{t}^{i}(x_{t}^{i},x)}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-r_{\bar{q}_{t}}\tilde{w}_{t}^{i}(x_{t}^{i},x)\right)dt.$ (88)
@@
$H(\bar{q}_{1},\bar{p}_{1})\leq H(\bar{q}_{0})-\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\tilde{w}_{t}^{i}(x^{i},x_{t})dt+D_{KL}(\bar{q}_{0}\|\bar{p}_{0})+\int_{0}^{1}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})$ (98)
$\sum_{x_{t}}\bar{q}_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\left(\frac{\bar{q}_{t}(x|x_{0},x_{1})}{\bar{q}_{t}(x_{t}|x_{0},x_{1})}\tilde{w}_{t}^{i}(x_{t}^{i},x)\log{\frac{\tilde{w}_{t}^{i}(x_{t}^{i},x)}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})\right)dt.$ (99)

### v4

@@ 5.5 Scaling Properties of Minibatch-OT
We examine the computational overhead introduced by minibatch OT training. While OT computation is vocabulary-size independent, its requirements increase with batch size. Table 4 compares the time for 1000 Sinkhorn couplings (CPU and GPU) against 1000 flow updates without OT, at fixed sequence length $L=128$, where we use the same batch size for OT as for flow training. OT adds only 3.4% overhead in our experiments. Larger batch sizes (1000-4096), as used by realistic LLMs such as GPT-3, PaLM 540B, LLaMA 3 405B, maintain the overhead around 12%, and require GPU acceleration for sizes of 2048-4096.
@@
Pure flow 54.7 63.6 104.9 173.0 367.8 634.4 $1129^{E}$ $2010^{E}$
Sequence length does not have a negative impact on computational scaling. Normally, the primary overhead stems from the solver (e.g. Sinkhorn) operation, which processes pre-computed pairwise sequence distances. Consequently, sequence length does not affect solver’s computational cost. We tested the overall role of the length empirically by increasing the sequence length 8 times (from $L=128$ to $L=1024$), which yielded only a 4.6x increase in OT computation time (from 12,57 to 57,99 seconds per thousand minibatches). This scaling behavior has important implications for training efficiency: At $L=128$, OT adds 3.4% to the total training time. At $L=1024$, this overhead should drop below 1.9% because flow-only training time scales at best roughly linearly (up to quadratically if attention dominates) with sequence length, whereas OT in our experiments scaled more slowly. Consequently, the relative cost of mini-batch OT is not expected to increase with sequence length.
In addition, minibatch OT scales favorably with embedding dimension $D$: OT computation scales linearly in $D$, while transformer computation scales quadratically. Consequently, as embedding dimensions grow, the relative overhead of our approach decreases compared to the flow optimization cost. Furthermore, minibatch-OT is agnostic to the number of transformer layers, meaning the relative overhead decreases even further as models scale.
@@ 6 Related Work
Diffusion-based models have proven highly effective in capturing the structure of continuous data distributions, leading to significant advancements in generative modeling (Song et al., 2020a, 2021c; Kingma et al., 2021; Nichol and Dhariwal, 2021; Saharia et al., 2022; Ramesh et al., 2022). Given their success in image, video and audio synthesis, researchers have explored their applicability to language modeling (Chen et al., 2023; Gulrajani and Hashimoto, 2024; Li et al., 2022; Dieleman et al., 2022; Strudel et al., 2022; Gong et al., 2022; Mahabadi et al., 2023).
An alternative paradigm for discrete data, particularly in NLP, is discrete diffusion. Introduced by Hoogeboom et al. (2021); Austin et al. (2021) and extended to continuous-time settings (Campbell et al., 2022; Lou et al., 2024), these models offer a structured approach to learning categorical distributions. Training typically uses the variational lower bound or cross-entropy loss, similar to continuous diffusion (Dieleman et al., 2022). To expand the design space of discrete diffusion, Campbell et al. (2024); Gat et al. (2024) introduce discrete flow matching, notably avoiding conditional score ratio calculations during training and thus bypassing matrix exponential computation. Instead of focusing on a path-length-oriented objective, Shaul et al. (2025) define a kinetic energy OT objective, derive the optimum for specific DFM classes, and independently obtain a bound similar to ours from an ELBO perspective. While in this work we focus on pure DFM models, Arriola et al. (2025) introduce block diffusion language models that interpolate between discrete denoising diffusion and autoregressive models. Regarding scaling, Nie et al. (2025) train masked diffusion models up to 1.1B parameters to systematically evaluate against comparable or larger ARMs. Their 1.1B MDM outperforms the 1.1B TinyLlama trained on the same data across four of eight zero-shot benchmarks.
@@ 7 Conclusion
We developed a weighted path-length dynamic OT objective for DFM that minimizes dissimilarity-weighted jumps between states, derived its Kantorovich formulation establishing a categorical Benamou-Brenier-type theorem. We extended two discrete diffusion bounds to DFM, enabling comparisons with autoregressive and discrete diffusion models. Experiments show minibatch OT reduces inference steps up to 32-fold (1024 to 32) while maintaining generative perplexity on GPT2-scale models. Our multimask flow surpasses masked DFM in generative perplexity without sacrificing diversity, with further gains under OT.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Proofs
is clearly $s(x_{1}^{i},x_{0}^{i})$ when $x_{0}^{i}\neq x_{1}^{i}$ and zero otherwise (therefore still $s(x_{1}^{i},x_{0}^{i})$ as $s(a,a)=0$). Thus, we have
@@
$D_{KL}(\hat{q},\hat{p})=\sum_{x_{0:K\epsilon}}\hat{q}(x_{0:K\epsilon})\log{\frac{\hat{q}(x_{0:K\epsilon})}{\hat{p}(x_{0:K\epsilon})}}=\sum_{x_{0:K\epsilon}}\hat{q}(x_{0:K\epsilon})\log{\bigg(\frac{\hat{q}(x_{0})}{\hat{p}(x_{0})}\prod_{k=1}^{K}}\frac{\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon},...,x_{0})}{\hat{p}(x_{k\epsilon}|x_{(k-1)\epsilon},...,x_{0})}\bigg)$ (37)
@@
First we will simplify notation and write $t_{k}=(k-1)\epsilon$, as well as $\hat{q}(x_{k\epsilon}=x|x_{(k-1)\epsilon}=z)=\hat{q}_{t_{k}+\epsilon|t_{k}}(x|z)$, where $z$ and $x$ are states. Therefore Expression (42) becomes
@@
$\sum_{\begin{subarray}{c}x\\ d_{H}(x,z)=0\end{subarray}}\hat{q}_{t_{k}+\epsilon|t_{k}}(x|z)\log{\frac{\hat{q}_{t_{k}+\epsilon|t_{k}}(x|z)}{\hat{p}_{t_{k}+\epsilon|t_{k}}(x|z)}}+\sum_{\begin{subarray}{c}x\\ d_{H}(x,z)=1\end{subarray}}\hat{q}_{t_{k}+\epsilon|t_{k}}(x|z)\log{\frac{\hat{q}_{t_{k}+\epsilon|t_{k}}(x|z)}{\hat{p}_{t_{k}+\epsilon|t_{k}}(x|z)}}$
@@
$\sum_{\begin{subarray}{c}x\\ d_{H}(x,z)=1\end{subarray}}\hat{q}_{t_{k}+\epsilon|t_{k}}(x|z)\log{\frac{\hat{q}_{t_{k}+\epsilon|t_{k}}(x|z)}{\hat{p}_{t_{k}+\epsilon|t_{k}}(x|z)}}=\sum_{j=1}^{L}\sum_{x^{j}\neq z^{j}}\epsilon w_{t_{k}}^{j}(x^{j},z)\log{\frac{w_{t_{k}}^{j}(x^{j},z)+O(\epsilon)}{v_{t_{k}}^{j}(x^{j},z)+O(\epsilon)}}+O(\epsilon^{2}).$ (48)
@@
$D_{KL}(\bar{q},\bar{p})=D_{KL}(\bar{q}_{0}\|\bar{p}_{0})+\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\left(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-w_{t}^{i}(x^{i},x_{t})\right)dt.$ (62)
@@
$D_{KL}(\bar{q},\bar{p})=D_{KL}(\bar{q}_{1}\|\bar{p}_{1})+\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\left(\tilde{w}_{t}^{i}(x^{i},x_{t})\log{\frac{\tilde{w}_{t}^{i}(x^{i},x_{t})}{\tilde{v}_{t}^{i}(x^{i},x_{t})}}+\tilde{v}_{t}^{i}(x^{i},x_{t})-\tilde{w}_{t}^{i}(x^{i},x_{t})\right)dt$ (63)
@@
$D_{KL}(\bar{q}_{1}\|\bar{p}_{1})=D_{KL}(\bar{q}_{0}\|\bar{p}_{0})+\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-w_{t}^{i}(x^{i},x_{t})\bigg)dt$
@@
$D_{KL}(\bar{q},\bar{p})=D_{KL}(\bar{q_{1}},\bar{p_{1}})$
$\int_{0}^{1}\sum_{z}\bar{q}_{1-t}(z)\sum_{i=1}^{L}\sum_{x^{i}\neq z^{i}}\bigg(r_{\bar{q}_{1-t}}w_{1-t}^{i}(z^{i},x)\log{\frac{r_{\bar{q}_{1-t}}w_{1-t}^{i}(z^{i},x)}{r_{\bar{p}_{1-t}}v_{1-t}^{i}(z^{i},x)}}+r_{\bar{p}_{1-t}}v_{1-t}^{i}(z^{i},x)-r_{\bar{q}_{1-t}}w_{1-t}^{i}(z^{i},x)\bigg)dt.$
@@
$D_{KL}(\bar{q},\bar{p})=D_{KL}(\bar{q_{1}},\bar{p_{1}})+\int_{0}^{1}\sum_{z}\bar{q}_{\tau}(z)\sum_{i=1}^{L}\sum_{x^{i}\neq z^{i}}\left(r_{\bar{q}_{\tau}}w_{\tau}^{i}(z^{i},x)\log{\frac{r_{\bar{q}_{\tau}}w_{\tau}^{i}(z^{i},x)}{r_{\bar{p}_{\tau}}v_{\tau}^{i}(z^{i},x)}}+r_{\bar{p}_{\tau}}v_{\tau}^{i}(z^{i},x)-r_{\bar{q}_{\tau}}w_{\tau}^{i}(z^{i},x)\right)d\tau$ (70)
@@
We now set to prove that $D_{KL}(\bar{q}(x_{0})\|\bar{p}(x_{0}))\leq D_{KL}(\bar{q},\bar{p})$ and $D_{KL}(\bar{q}(x_{1})\|\bar{p}(x_{1}))\leq D_{KL}(\bar{q},\bar{p})$. Since in Equation (65), the term
@@
$D_{KL}(\bar{q_{1}}\|\bar{p_{1}})\leq D_{KL}(\bar{q_{0}}\|\bar{p_{0}})+\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\left(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-w_{t}^{i}(x^{i},x_{t})\right)dt.$ (72)
@@
$D_{KL}(\bar{q}_{1}\|\bar{p}_{1})\leq D_{KL}(\bar{q}_{0}\|\bar{p}_{0})+\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\left(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-w_{t}^{i}(x^{i},x_{t})\right)dt$ (76)
@@
$\frac{\partial H(\bar{q}_{t})}{\partial t}=-\frac{\partial}{\partial t}\sum_{x_{t}}\bar{q}_{t}(x_{t})\log{\bar{q}_{t}(x_{t})}=-\sum_{x_{t}}\frac{\partial\bar{q}_{t}(x_{t})}{\partial t}\left(\log{\bar{q}_{t}(x_{t})}+1\right)=\sum_{x_{t}}\frac{\partial\bar{q}_{t}(x_{t})}{\partial t}\left(\log{\frac{1}{\bar{q}_{t}(x_{t})}}-1\right).$ (78)
From the Continuity Equation (Gat et al., 2024),
@@
$\frac{\partial H(\bar{q}_{t})}{\partial t}=\sum_{x_{t}}\frac{\partial\bar{q}_{t}(x_{t})}{\partial t}\left(\log{\frac{1}{\bar{q}_{t}(x_{t})}}-1\right)=\sum_{x_{t}}\sum_{x}\bar{q}_{t}(x)\sum_{i=1}^{L}\delta_{x}(x_{t}^{-i})w_{t}^{i}(x_{t}^{i},x)\left(\log{\frac{1}{\bar{q}_{t}(x_{t})}}-1\right).$ (80)
We note that
$\frac{\partial H(\bar{q}_{t})}{\partial t}=\sum_{x_{t}}\sum_{x}\bar{q}_{t}(x)\sum_{i=1}^{L}\delta_{x}(x_{t}^{-i})w_{t}^{i}(x_{t}^{i},x)\left(\log{\frac{1}{\bar{q}_{t}(x_{t})}}-1\right)$ (81)
$\sum_{x_{t}}\sum_{x}\bar{q}_{t}(x)\sum_{i=1}^{L}\delta_{x}(x_{t}^{-i})w_{t}^{i}(x_{t}^{i},x)\left(\log{\frac{\bar{q}_{t}(x)}{\bar{q}_{t}(x_{t})}}-1\right)-\sum_{x_{t}}\sum_{x}\bar{q}_{t}(x)\sum_{i=1}^{L}\delta_{x}(x_{t}^{-i})w_{t}^{i}(x_{t}^{i},x)\log{\bar{q}_{t}(x)}.$ (82)
The last term is 0. Indeed,
$\sum_{x}\bar{q}_{t}(x)\sum_{i=1}^{L}\sum_{x_{t}}\left(\delta_{x}(x_{t}^{-i})w_{t}^{i}(x_{t}^{i},x)\right)\log{\bar{q}_{t}(x)}=\sum_{x}\bar{q}_{t}(x)\sum_{i=1}^{L}0\log{\bar{q}_{t}(x)}=0,$ (83)
thus
$\frac{\partial H(\bar{q}_{t})}{\partial t}=\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}}w_{t}^{i}(x_{t}^{i},x)\frac{\bar{q}_{t}(x)}{\bar{q}_{t}(x_{t})}\left(\log{\frac{\bar{q}_{t}(x)}{\bar{q}_{t}(x_{t})}}-1\right).$ (84)
@@
$D_{KL}(\bar{q}_{1}\|\bar{p}_{1})\leq D_{KL}(\bar{q}_{0}\|\bar{p}_{0})+\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\left(r_{\bar{q}_{t}}\tilde{w}_{t}^{i}(x_{t}^{i},x)\log{\frac{r_{\bar{q}_{t}}\tilde{w}_{t}^{i}(x_{t}^{i},x)}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-r_{\bar{q}_{t}}\tilde{w}_{t}^{i}(x_{t}^{i},x)\right)dt.$ (86)
@@
$H(\bar{q}_{1},\bar{p}_{1})\leq H(\bar{q}_{0})-\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\tilde{w}_{t}^{i}(x^{i},x_{t})dt+D_{KL}(\bar{q}_{0}\|\bar{p}_{0})$ (95)
$\int_{0}^{1}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})\sum_{x_{t}}\bar{q}_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\left(\frac{\bar{q}_{t}(x|x_{0},x_{1})}{\bar{q}_{t}(x_{t}|x_{0},x_{1})}\tilde{w}_{t}^{i}(x_{t}^{i},x)\log{\frac{\tilde{w}_{t}^{i}(x_{t}^{i},x)}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})\right)dt.$ (96)